Question 9

Sheridan Corporation manufactures a single product. The standard cost per unit of product is shown below.

Direct materials—1 pound plastic at $7.00 per pound $ 7.00
Direct labor—1.5 hours at $11.10 per hour 16.65
Variable manufacturing overhead 9.00
Fixed manufacturing overhead 9.00
Total standard cost per unit $41.65


The predetermined manufacturing overhead rate is $12.00 per direct labor hour ($18.00 ÷ 1.5). It was computed from a master manufacturing overhead budget based on normal production of 9,000 direct labor hours (6,000 units) for the month. The master budget showed total variable costs of $54,000 ($6.00 per hour) and total fixed overhead costs of $54,000 ($6.00 per hour). Actual costs for October in producing 3,500 units were as follows.

Direct materials (3,660 pounds) $ 25,986
Direct labor (5,140 hours) 59,110
Variable overhead 44,098
Fixed overhead 21,202
    Total manufacturing costs $150,396


The purchasing department buys the quantities of raw materials that are expected to be used in production each month. Raw materials inventories, therefore, can be ignored.

(a)

Compute all of the materials and labor variances.

Total materials variance $

Neither favorable nor unfavorableFavorableUnfavorable

Materials price variance $

UnfavorableNeither favorable nor unfavorableFavorable

Materials quantity variance $

Neither favorable nor unfavorableFavorableUnfavorable

Total labor variance $

UnfavorableNeither favorable nor unfavorableFavorable

Labor price variance $

FavorableNeither favorable nor unfavorableUnfavorable

Labor quantity variance $

FavorableUnfavorableNeither favorable nor unfavorable


(b)

Compute the total overhead variance.

Total overhead variance $

Neither favorable nor unfavorableUnfavorableFavorable
Answer #1

Solution a:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
3660 $7.10 $25,986.00 3660 $7.00 $25,620.00 3500 $7.00 $24,500.00
$366.00 Unfavorable $1,120.00 Unfavorable
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $366.00 Unfavorable
Direct material quantity variance $1,120.00 Unfavorable
Total direct material variance $1,486.00 Unfavorable
Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
5140 $11.50 $59,110.00 5140 $11.10 $57,054.00 5250 $11.10 $58,275.00
$2,056.00 Unfavorable $1,221.00 Favorable
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $2,056.00 Unfavorable
Direct Labor Efficiency variance $1,221.00 Favorable
Total direct labor variance $835.00 Unfavorable



Solution b:

Total overhead variance = Total overhead applied - Actual overhead

= (3500*1.5*$12) - ($44,098 + $21,202)

= $63,000 - $65,300 = $2,300 U